if 261-h=67,what is the value of h
Answer by ptaylor(1894)


261-h=67 First, subtract 261 from both sides of the equation
261-261-h=67-261 Next, collect like terms
-h=-194 Now we will multiply both sides by -1
h=194--------------ans
Hope this helps----ptaylor